### v1

# Scalable Binary CUR Low-Rank Approximation Algorithm

Scalable Binary CUR Low-Rank Approximation Algorithm
Bowen Su
Department of Math
Michigan State University
I Introduction
Large-scale matrices from applications such as social networks [21] and genomics [2] present significant challenges in modern data analysis. These matrices often exhibit low-rank structures, which can be exploited through matrix factorization techniques to derive compact and informative representations. Classical methods for low-rank matrix approximation, including Singular Value Decomposition (SVD) [15], rank-revealing QR [20], and rank-revealing LU decomposition [25], are well-established and provide reliable solutions. However, their computational costs scale rapidly with matrix size. Specifically, SVD has a complexity of $\mathcal{O}(n^{2}m)$ for a matrix $A\in\mathbb{R}^{n\times m}$ with $n\geq m$, while rank-revealing QR and LU decomposition require $\mathcal{O}(nm^{2})$. Such high computational demands render these methods impractical for applications involving extremely large datasets, creating significant challenges in terms of scalability and efficiency for large-scale data analysis. To address these limitations, CUR Low-Rank Approximaation [11, 7] has emerged as a potential alternative, offering efficient approximation of matrices with inherent low-rank structures.
I-A CUR Low-Rank Approximaation
Let $A\in\mathbb{R}^{m\times n}$ be a given matrix, and consider two sets of indices $I=\{i_{1},\cdots,i_{k}\}$ and $J=\{j_{1},\cdots,j_{k}\}$, corresponding to selected rows and columns, respectively. Define $A_{I,:}$ as the subset of rows in $A$ indexed by $I$, and $A_{:,J}$ as the subset of columns indexed by $J$. Suppose the submatrix $A_{I,J}$, formed by the intersection of rows $I$ and columns $J$, is nonsingular. In this case, the approximation
$A\approx A_{:,J}A_{I,J}^{\dagger}A_{I,:}$
is referred to as a CUR approximiation of $A$. For a visual representation, see Figure 1.
Figure 1: CUR Approximation
There has been extensive research on CUR Low-Rank Approximations for matrices. For further details, refer to [7, 14, 19, 22, 30, 32, 31] and the references therein. There are different strategies to obtain such an approximations or decomposition in the literature. These include volume optimization [9, 16, 23, 17, 1, 28, 18, 34], rank and spectrum-revealing factorization [20, 3], probabilistic sampling [32, 31, 33, 7, 12, 22], and empirical interpolation [8, 13, 30]. This paper focuses on the first category within the framework of parallel computing.
I-B Maximal Volume Optimization
Definition 1
Given a square matrix $A\in\mathbb{R}^{r\times r},$ the volume of it is defined as absolute value of its determinant, i.e.,
$\operatorname{vol}(A)=|det(A)|.$
The CUR approximation with $r$ selected columns and rows is usually formulated as the following optimization problem [9, 16, 23, 17, 1, 28, 18, 34]:
$\operatorname*{argmax}_{\begin{subarray}{c}|\hat{I}|=|\hat{J}|=r\end{subarray}}\operatorname{vol}(A(\hat{I},\hat{J})).$
The aforementioned optimization problem has numerous practical applications such as identifying the optimal nodes for polynomial interpolation over a given domain on appropriately discretized meshes [29] and preconditioning the iterative solution of linear least-squares problems [4].
Theorem 1 (cf. [17])
Let $A$ be a matrix of size $m\times n$ and $r$ be an integer satisfying $1\leq r<\min\{m,n\})$. Suppose that $I\subset\{1,\cdots,m\}$ with #$\#(I)=r$ and $J\subset\{1,\cdots,n\}$ with #$\#(J)=r$ such that $A_{I,J}$ has the maximal volume among all $r\times r$ submatrices of $A$. Then, the Chebyshev norm of the residual matrix satisfies the following estimate:
$\|A-A_{r}\|_{C}\leq(r+1)\sigma_{r+1}(A),$ (1)
where $A_{r}=A_{:,J}A_{I,J}^{\dagger}A_{I,:}$ is the $r$-cross approximation of $A$, $\sigma_{r+1}(A)$ is the $(r+1)^{th}$ singular value of $A$.
Recently, under the assumption in Theorem 1, Allen et.al [1] improve the error estimate bounds to the following the results.
Theorem 2 (cf. [1])
Under the assumption in Theorem 1, there is following estimate:
$\|A-A_{r}\|_{C}\leq\frac{(r+1)\sigma_{r+1}(A)}{\sqrt{1+\sum\limits_{k=1}^{r}\frac{\sigma_{r+1}^{2}(A)}{\sigma_{k}^{2}(A)}}}$
Theorem 1 and Theorem 2 tells us that the volume maximization can lead to a quasi-best low-rank approximation. However, identifying maximum-volume submatrices presents considerable computational challenges, as the problem is proven to be NP-hard [5]. This complexity has driven the development and adoption of greedy method-adaptive cross approximation-to provide practical solutions [6]. At its core, this method is highly connected to incomplete LU decomposition of the matrix [10]. We firstly introduce CUR Low-Rank Approximaation via Adaptive Cross Approximation Algorithm (ACA) [6] as follows.
Algorithm 1 Scalable Binary CUR Low-Rank Approximation Algorithm
1:Input: Matrix $A\in\mathbb{R}^{m\times n}$, target rank $r$.
2:Initialize $R_{0}:=A$, $I:=\{\},J:=\{\}.$
3:for $k=0,\ldots,m-1$ do
4:     $(i_{k+1},j_{k+1}):=\operatorname*{argmax}_{i,j}|R_{k}(i,j)|.$
5:     $I\leftarrow I\cup\{i_{k+1}\},J\leftarrow J\cup\{j_{k+1}\}.$
6:     $R_{k+1}:=R_{k}-\frac{1}{R_{k}(i_{k+1},j_{k+1})}R_{k}(:,j_{k+1})R_{k}(i_{k+1},:).$
7:end for
8:Output: $C=A(:,J),U=A(I,J),R=A(I,:).$
Each iteration in Algorithm 1 identifies a single pivot by searching for the largest absolute value in the residual matrix. Although local maxima searches can be conducted in parallel, a global reduction is necessary to determine the unique pivot element, thus creating an unavoidable sequential bottleneck. Once the pivot is found, the algorithm performs a rank-1 update on the residual, and this updated matrix in turn dictates the location and value of the next pivot. Consequently, there is a strict iteration-by-iteration data dependency, preventing the selection of multiple pivots simultaneously. In this paper, we propose the Scalable Binary CUR Low-Rank Approximation Algorithm. This framework is specifically designed to enable scalable CUR approximations for large-scale matrices.
II Scalable Binary CUR Low-Rank Approximaation Algorithm
II-A Overview
The proposed Scalable Binary CUR Low-Rank Approximation Algorithm efficiently approximates a large-scale matrix $A\in\mathbb{R}^{n\times m}$ by decomposing it into three smaller matrices, $C$, $R$, and $U$. First, a binary parallel selection process, as detailed in Algorithm 3, identifies representative subsets of rows and columns from $A$. The matrix $C$ is constructed by extracting the columns of $A$ corresponding to the selected indices, while $R$ is formed by extracting the rows based on these indices. Finally, $U$ is computed as the pseudoinverse of the intersection submatrix $A(\text{rows},\text{cols})$. Together, these three matrices yield an efficient low-rank approximation of $A$.
Algorithm 2 Scalable Binary CUR Low-Rank Approximaation Algorithm
1:Input: Matrix $A\in\mathbb{R}^{n\times m}$, selected row count $r$, selected column count $c$
2:// Perform the following steps in parallel
3:Use Algorithm 3 to determine the set of row indices $I$.
4:Use Algorithm 3 to determine the set of column indices $J$.
5:Output: $C=A(:,J),U=A(I,J),R=A(I,:).$
Algorithm 3 Blockwise Adaptive Cross Approximation
1:Input: Matrix $A\in\mathbb{R}^{n\times m}$, number of selected indices $r$, number of blocks $b$, axis (0 for rows, 1 for columns)
2:Output: Selected indices $I$
3:$I=\{\}$triangleright Initialize selected indices
4:if axis = 0 then
5:     Divide $A$ into $b$ blocks along rows: $A_{\text{block}}^{k}\in\mathbb{R}^{n_{k}\times m}$
6:     dimBlockSize $\leftarrow n_{k}$, dimSize $\leftarrow n$
7:else
8:     Divide $A$ into $b$ blocks along columns: $A_{\text{block}}^{k}\in\mathbb{R}^{n\times m_{k}}$
9:     dimBlockSize $\leftarrow m_{k}$, dimSize $\leftarrow m$
10:end if
11:for $i=0$ to $r-1$ do $\triangleright$ Select $r$ indices
12:     localMaxNorm, localMaxIdx $\leftarrow-1,-1$
13:     // Blockwise parallel selection of rows/columns with maximal norm
14:     for $i=0$ to $b\cdot\texttt{dimBlockSize}-1$ (parallel) do
15:         $k\leftarrow\lfloor i/\texttt{dimBlockSize}\rfloor$
16:         $j\leftarrow i\bmod\texttt{dimBlockSize}$
17:         $\texttt{globalIdx}\leftarrow i$
18:         if axis = 0 then
19:              $\texttt{localNorm}\leftarrow\operatorname{norm}(A_{\text{block}}^{k}(j,:))$
20:         else
21:              $\texttt{localNorm}\leftarrow\operatorname{norm}(A_{\text{block}}^{k}(:,j))$
22:         end if
23:         if localNorm $>$ localMaxNorm then
24:              localMaxNorm $\leftarrow$ localNorm
25:              localMaxIdx $\leftarrow$ globalIdx
26:         end if
27:     end for
28:     for $k=0$ to $b-1$ (parallel) do
29:         if localMaxNorm $>$ maxNorm then
30:              maxNorm $\leftarrow$ localMaxNorm
31:              maxIdx $\leftarrow$ localMaxIdx
32:         end if
33:     end for
34:     $k_{\text{block}}\leftarrow\texttt{maxIdx}\div\texttt{dimBlockSize}$
35:     $j_{\text{local}}\leftarrow\texttt{maxIdx}\bmod\texttt{dimBlockSize}$
36:     if axis = 0 then
37:         sharedMaxVec $\leftarrow A_{\text{block}}^{k}(j_{\text{local}},:)$
38:     else
39:         sharedMaxVec $\leftarrow A_{\text{block}}^{k}(:,j_{\text{local}})$
40:     end if
41:     // Parallel Blockwise update using shared max vector
42:     for $k=0$ to $b-1$ (parallel) do
43:         if axis = 0 then
44:
$A_{\text{block}}^{k}\leftarrow A_{\text{block}}^{k}-\frac{A_{\text{block}}^{k}\cdot\texttt{sharedMaxVec}^{\top}}{\texttt{maxNorm}}\cdot\texttt{sharedMaxVec}$
45:         else
46:
$A_{\text{block}}^{k}\leftarrow A_{\text{block}}^{k}-\texttt{sharedMaxVec}\cdot\left(\frac{\texttt{sharedMaxVec}^{\top}\cdot A_{\text{block}}^{k}}{\texttt{maxNorm}}\right)$
47:         end if
48:     end for
49:     $I(i)\leftarrow\texttt{maxIdx}$
50:end for
51:return $I$
II-B Blockwise Adaptive Cross Algorithm
The Blockwise Adaptive Cross Approximation algorithm is designed to efficiently select a subset of $r$ representative indices from a large matrix $A\in\mathbb{R}^{n\times m}$ by iteratively extracting the most significant row or column vectors. The algorithm begins by partitioning the matrix into $b$ blocks along a specified axis, determined by the parameter axis (where 0 indicates a row-wise partition and 1 indicates a column-wise partition). When axis is set to 0, $A$ is divided into blocks $A_{\text{block}}^{k}\in\mathbb{R}^{n_{k}\times m}$ along the rows; if axis is set to 1, the matrix is partitioned into blocks $A_{\text{block}}^{k}\in\mathbb{R}^{n\times m_{k}}$ along the columns. In this context, the variable dimBlockSize refers to the number of rows or columns in each block, while dimSize represents the total number of rows $n$ or columns $m$ in the matrix.
An empty global index set $I$ is then initialized to store the indices of the selected rows or columns. The algorithm proceeds over $r$ iterations, each intended to identify one key index. In each iteration, a parallel search is conducted within each block to compute the Euclidean norm of every candidate vector (row or column, depending on the specified axis). Each block determines its local maximum norm and the corresponding local index. These local maxima are subsequently combined through a reduction step that identifies the global maximum norm and its associated index, denoted as maxIdx.
Once the vector corresponding to maxIdx is identified, it is extracted from its block and referred to as sharedMaxVec. This vector is then employed to perform a rank-1 update on every block of the matrix. Specifically, if the operation is row-wise (i.e., axis = 0), the update involves subtracting from each block the product of the block and the transpose of sharedMaxVec divided by the squared norm of sharedMaxVec, multiplied by sharedMaxVec itself. Conversely, if the operation is column-wise (i.e., axis = 1), the subtraction is carried out by multiplying sharedMaxVec with the appropriate normalized factor computed from the block. This subtraction effectively removes the dominant component captured by the selected vector, ensuring that subsequent iterations focus on extracting new, non-redundant information.
The complete pseudo code is provided in Algorithm 3 for detailed reference.
The complexity of the Block Adaptive Cross Algorithm is determined by the operations performed in each iteration. Computing row norms for $n_{k}$ rows within a block requires $\mathcal{O}(n_{k}m)$ operations, and with $b$ blocks, this totals $\mathcal{O}\left(\frac{nm}{b}\right)$ per iteration. Updating the matrix, which includes projection, also requires $\mathcal{O}\left(\frac{nm}{b}\right)$ operations per iteration.
For $r$ iterations, the total complexity of the algorithm is: $\mathcal{O}\left(r\cdot\frac{nm}{b}\right).$ This represents a significant improvement over the sequential Adaptive Cross Approximation algorithm’s complexity of $O(r\cdot nm)$, particularly for large matrices and high levels of parallelism.
III Numerical Experiments
All experiments are conducted in a computing environment powered by an AMD EPYC 7H12 64-Core Processor running at 2.6 GHz. For matrix operations, we utilized Armiddol[26, 27], a high-performance matrix library designed to efficiently handle large-scale linear algebra computations.
III-A Approximation Error Study of Algorithm 2
In this section, we evaluate the performance of Algorithm 2 by testing its approximation capabilities on two types of data: Hilbert matrices and synthetic low-rank matrices. The key metric used for evaluation is the Frobenius norm relative error, defined as:
Relative $\text{Relative Error}=\frac{\|A-CUR\|_{F}}{\|A\|_{F}},$
where $C$, $U$, and $R$ are the matrices resulting from the CUR Low-Rank Approximaation of $A$. This metric provides a quantitative assessment of the quality of the approximation and is consistently applied across all experiments.
Experiments on Hilbert Matrices
Hilbert matrices are well-known for their ill-conditioned nature and high sensitivity to numerical approximations [24]. The experiments assess the approximation error of Algorithm 2 applied to Hilbert matrices of varying dimensions ($256\times 256$, $512\times 512$, and $1024\times 1024$). For each matrix size, the number of selected rows and columns varied from $1$ to $20$. Figure 2 shows the relationship between the number of selected rows and columns and the approximation error.
Figure 2: Approximation error of Algorithm 2 for Hilbert matrices. The x-axis represents the number of selected rows/columns, and the y-axis shows the Frobenius norm relative error.
The experimental results, illustrated in Figure 2, demonstrate the efficacy of Algorithm 2 in approximating matrices of varying sizes ($256\times 256$, $512\times 512$, and $1024\times 1024$) by examining the relationship between the number of selected rows/columns and the Frobenius norm relative error. Figure 2 reflects the algorithm’s ability to effectively capture the dominant features of the matrix.
Experiments on Synthetic Low-Rank Matrices
To further validate the algorithm, we conducted tests on synthetic low-rank matrices with controlled rank structures. These matrices were constructed by generating two matrices, $A$ and $B$, of sizes $n\times r$ and $r\times n$, respectively, with entries defined as:
$A(i,j)=\frac{1}{i+j},\quad B(i,j)=\frac{1}{i+j},$
where $i$ and $j$ denote the row and column indices. The synthetic low-rank matrix $H$ is generated by $H=A\cdot B.$ The experiments are performed on the generated matrix with sizes $n=256,512,1024$ and rank value $r=5,10,15,20$.
(a) Rank = 5
(b) Rank = 10
(c) Rank = 15
(d) Rank = 20
Figure 3: Relative Errors for Different Ranks
The results reveal that the Algorithm 2 achieves near-perfect approximation for low-rank matrices even with a small number of selected rows and columns. The relative error decreases sharply as $r$ increases and stabilizes at near-zero levels, reflecting the algorithm’s effectiveness in reconstructing matrices with well-defined rank structures.
III-B Scalability Analysis of Algorithm 2
In this section, we conduct a scalability analysis of Algorithm 2. The experiments focus on matrices of size $16384\times 16384$. We choose Algorithm 1 as sequential benchmark algorithm
Figure 4: Caption
Figure 4 displays the relationship between the number of processes and the total execution time for the Algorithm 2. As the number of processes increases from 2 to 64, the execution time exhibits a decreasing trend, demonstrating the expected performance improvement with parallelization. Notably, the execution time reduces from 12.3733 seconds for two process to 1.02225 seconds when employing 64 processes. However, the performance gain is not strictly linear; the curve gradually flattens as the number of processes increases beyond 16, suggesting diminishing returns due to parallel overhead and potential communication costs. A red dashed line is plotted to indicate the sequential baseline time of 7.4271 seconds, which serves as a reference point. The data highlights that parallel execution with more than 4 processes consistently outperforms the sequential baseline. These results suggest that parallelization is effective in accelerating MaxVol computations, but the efficiency gain saturates at higher process counts, likely due to inter-process communication and synchronization overhead.
IV Conclusion
In conclusion, we have presented a Scalable Binary CUR Low-Rank Approximation Algorithm. This approach is designed to overcome the scalability challenges inherent in the CUR decomposition of large-scale matrices in a deterministic way. By leveraging the parallel processing capabilities of modern multi-core architectures, our approach has shown modest improvements in computational efficiency. Future work will focus on refining and optimizing the algorithm to achieve more reliable outcomes in practical data processing scenarios.

### v2

@@
 Bowen Su
-Department of Math
-Michigan State University
 I Introduction
@@
 I-A CUR Low-Rank Approximaation
-Let $A\in\mathbb{R}^{m\times n}$ be a given matrix, and consider two sets of indices $I=\{i_{1},\cdots,i_{k}\}$ and $J=\{j_{1},\cdots,j_{k}\}$, corresponding to selected rows and columns, respectively.
+Let $A\in\mathbb{R}^{m\times n}$ be a given matrix, and consider two sets of indices $I\subset[m]$ and $J\subset[n]$, corresponding to selected rows and columns, respectively.
 Define $A_{I,:}$ as the subset of rows in $A$ indexed by $I$, and $A_{:,J}$ as the subset of columns indexed by $J$.
@@
 In this case, the approximation
-$A\approx A_{:,J}A_{I,J}^{\dagger}A_{I,:}$
-is referred to as a CUR approximiation of $A$. For a visual representation, see Figure 1.
+$A\approx CUR,$
+where $C=A(:,J),U=A^{\dagger}(I,J),R=A(I,:),$ is referred to as a CUR approximiation of $A$. For a visual representation, see Figure 1.
 Figure 1: CUR Approximation
 There has been extensive research on CUR Low-Rank Approximations for matrices.
-For further details, refer to [7, 14, 19, 22, 30, 32, 31] and the references therein.
-There are different strategies to obtain such an approximations or decomposition in the literature.
-These include volume optimization [9, 16, 23, 17, 1, 28, 18, 34], rank and spectrum-revealing factorization [20, 3], probabilistic sampling [32, 31, 33, 7, 12, 22], and empirical interpolation [8, 13, 30].
-This paper focuses on the first category within the framework of parallel computing.
+For readers interested in further details and additional references, please see the following works [1, 2, 4, 5, 9, 10, 11, 12, 13, 14, 15, 17, 19, 20, 21, 22, 23, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 36, 40, 41, 46, 47, 48, 49, 50, 51, 52, 53] and the references cited therein.
+There are different strategies to obtain such CUR approximations or decomposition.
+This paper explore maximal volume optimization strategy within the framework of parallel computing.
 I-B Maximal Volume Optimization
@@
 $\operatorname{vol}(A)=|det(A)|.$
-The CUR approximation with $r$ selected columns and rows is usually formulated as the following optimization problem [9, 16, 23, 17, 1, 28, 18, 34]:
+The CUR approximation with $r$ selected columns and rows is usually formulated as the following optimization problem [2, 15, 25, 26, 27, 37, 44, 53]
 $\operatorname*{argmax}_{\begin{subarray}{c}|\hat{I}|=|\hat{J}|=r\end{subarray}}\operatorname{vol}(A(\hat{I},\hat{J})).$
@@
 In this paper, we propose the Scalable Binary CUR Low-Rank Approximation Algorithm.
-This framework is specifically designed to enable scalable CUR approximations for large-scale matrices.
+This framework is specifically designed to explore scalable CUR approximations for large-scale matrices by selecting rows and columns in parallel.
 II Scalable Binary CUR Low-Rank Approximaation Algorithm
@@
 In conclusion, we have presented a Scalable Binary CUR Low-Rank Approximation Algorithm.
-This approach is designed to overcome the scalability challenges inherent in the CUR decomposition of large-scale matrices in a deterministic way.
+This approach is designed to achieve CUR decomposition of large-scale matrices by deterministiclly selecting rows and columns in parallel.
 By leveraging the parallel processing capabilities of modern multi-core architectures, our approach has shown modest improvements in computational efficiency.
-Future work will focus on refining and optimizing the algorithm to achieve more reliable outcomes in practical data processing scenarios.